# Enhancing nanocrystal superlattice self-assembly near a metastable liquid binodal

### v1

@@ Acknowledgments
We thank M. Delor for help designing the DLS setup. This work was supported by the Office of Basic Energy Sciences (BES), US Department of Energy (DOE) (award no. DE-SC0019375). Use of the Stanford Synchrotron Radiation Lightsource, SLAC National Accelerator Laboratory, is supported by the DOE, Office of Science, Office of Basic Energy Sciences (contract no. DE-AC02-76SF00515). Use of beamline 7.3.3 at the Advanced Light Source, Lawrence Berkeley National Laboratory is supported by the DOE, Office of Science, Office of Basic Energy Sciences. C.P.N.T., V.R.K.W., and R.B.W. were supported by the NSF Graduate Research Fellowship. L.M.H. and J.A.T. acknowledge a National Defense Science and Engineering Graduate Fellowship. J.K.U. was supported by an Arnold O. Beckman Postdoctoral Fellowship in Chemical Sciences from the Arnold and Mabel Beckman Foundation. A.D. was supported by a Philomathia Graduate Student Fellowship from the Kavli Nanoscience Institute at UC Berkeley. D.T.L. was supported by an Alfred P. Sloan Research Fellowship. N.S.G. was supported by a David and Lucile Packard Foundation Fellowship for Science and Engineering and Camille and a Henry Dreyfus Teacher-Scholar Award.

### v2

@@ Acknowledgments
We thank M. Delor for help designing the DLS setup. This work was supported by the Office of Basic Energy Sciences (BES), US Department of Energy (DOE) (award no. DE-SC0019375). Use of the Stanford Synchrotron Radiation Lightsource, SLAC National Accelerator Laboratory, is supported by the DOE, Office of Science, Office of Basic Energy Sciences (contract no. DE-AC02-76SF00515). Use of beamline 7.3.3 at the Advanced Light Source, Lawrence Berkeley National Laboratory is supported by the DOE, Office of Science, Office of Basic Energy Sciences (contract no. DE-AC02-05CH11231). C.P.N.T., V.R.K.W., and R.B.W. were supported by the NSF Graduate Research Fellowship. L.M.H. and J.A.T. acknowledge a National Defense Science and Engineering Graduate Fellowship. J.K.U. was supported by an Arnold O. Beckman Postdoctoral Fellowship in Chemical Sciences from the Arnold and Mabel Beckman Foundation. A.D. and L.M.H. were supported by Philomathia Graduate Student Fellowships from the Kavli Energy NanoScience Institute at UC Berkeley. D.T.L. was supported by an Alfred P. Sloan Research Fellowship. N.S.G. was supported by a David and Lucile Packard Foundation Fellowship for Science and Engineering and Camille and a Henry Dreyfus Teacher-Scholar Award.